# Phantom Crossing and Oscillating Dark Energy with $F(R)$ Gravity

### v1

@@ Abstract
In this work, we shall consider how a dynamical oscillating and phantom crossing dark energy era can be realised in the context of $F(R)$ gravity. We approach the topic from a theoretical standpoint, considering all the conditions that may lead to a consistent phantom crossing behaviour and separately how the $F(R)$ gravity context may realise an oscillating dark energy era. Apart from our qualitative considerations, we study quantitatively two $F(R)$ gravity dark energy models, which are viable cosmologically and also exhibit simultaneously phantom crossing behaviour and oscillating dark energy. We consider these models by numerically solving the field equations using appropriate statefinder parameters engineered for dark energy studies. As we show, $F(R)$ provides a natural extension of Einstein’s general relativity, which can naturally realise a transition from a phantom era to a quintessential era, a feature supported by recent observational data, without resorting to phantom scalar fields to realise the phantom evolution.
preprint: KEK-TH-2734, KEK-Cosmo-0383
@@ I Introduction
The observation of the late-time acceleration of the Universe, known as dark energy (DE) era is one of the most unexpected features of our world. In the context of simple general relativity, the DE era can be generated by quintessence scalar fields, however, the latest observations of the Dark Energy Spectroscopic Instrument (DESI) DESI:2024mwx ; DESI:2025zgx indicate that the DE was in the phantom regime in the near past. This feature can be described in the context of general relativity only by using phantom scalar fields, which is a rather unattractive feature. Also and more importantly, the DESI data indicate a transition from a phantom to a quintessential regime for DE. This is questionable if it can be realised in the context of general relativity. There are many ways to produce results compatible with the DESI data, for example Odintsov:2024woi ; Dai:2020rfo ; He:2020zns ; Nakai:2020oit ; DiValentino:2020naf ; Agrawal:2019dlm ; Ye:2020btb ; Vagnozzi:2021tjv ; Desmond:2019ygn ; Hogas:2023pjz ; OColgain:2018czj ; Vagnozzi:2019ezj ; Krishnan:2020obg ; Colgain:2019joh ; Vagnozzi:2021gjh ; Lee:2022cyh ; Krishnan:2021dyb ; Ye:2021iwa ; Ye:2022afu ; Verde:2019ivm ; Menci:2024rbq ; Adil:2023ara ; Reeves:2022aoi ; Ferlito:2022mok ; Vagnozzi:2021quy ; DiValentino:2020evt ; Sabogal:2024yha ; Giare:2024smz ; DiValentino:2025sru ; Odintsov:2025kyw ; vanderWesthuizen:2025iam ; Paliathanasis:2025dcr ; Odintsov:2025jfq ; Kessler:2025kju ; Nojiri:2025jhh ; Ozulker:2025ehg and among these frameworks, $F(R)$ gravity Nojiri:2017ncd ; Capozziello:2011et ; Nojiri:2006ri ; Nojiri:2010wj ; Nojiri:2003ft ; Capozziello:2005ku ; Hwang:2001pu ; Song:2006ej ; Faulkner:2006ub ; Olmo:2006eh ; Sawicki:2007tf ; Faraoni:2007yn ; Carloni:2007yv ; Nojiri:2007as ; Deruelle:2007pt ; Appleby:2008tv ; Dunsby:2010wg ; Oikonomou:2025qub enjoys an elevated role. There are recent studies which discuss the ability of $F(R)$ gravity to be compatible with the DESI data Odintsov:2025jfq . Another more fascinating behaviour of DE based on the DESI data was demonstrated in Ref. Kessler:2025kju , which indicates that the DE might not be only dynamical and evolving from phantom to quintessence, but it also shows strong indications of an oscillating equation of state (EoS). In this work, we aim to investigate whether a phantom involving and oscillating DE may be realised by $F(R)$ gravity in general. We approach the problem theoretically, finding the conditions which may allow such a phenomenological behaviour to be produced, and also we present several phenomenological models from a theoretical point of view. In addition, we study numerically the DE era of some explicit $F(R)$ gravity models which can realise a viable Planck compatible Planck:2018vyg and in some cases DESI compatible late-time behaviour. We also demonstrate that these models produce an oscillating DE era and also allow for phantom to quintessence transitions.
@@
This paper is organised as follows: In Section II, we discuss the inverse-phantom processing in $F(R)$ gravity from a theoretical viewpoint. The analysis is continued in Section III, where we discuss the apparent phantom crossing. In Section IV, the theoretical study of several models that exhibit phantom crossing is considered. The oscillating Hubble rate is considered in Section V, while in Section VI, the effects of a scalaron as dark matter with curvature-dependent mass on the DE era are considered. In Section VII, we discuss the features of a realistic model, while in Section VIII, we numerically study two explicit examples of oscillating $F(R)$ gravity DE with phantom crossing. Finally, the conclusions follow at the end of the article.
@@ II Inverse-Phantom Crossing in $F(R)$ gravity
$\mathcal{S}=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}F(R)\,,$ (1)
@@
Let $w$ be the DE equation of state (EoS) parameter, that is, the ratio of the pressure of the DE fluid with respect to the energy density of the DE fluid. If the recent DESI observations DESI:2024mwx ; DESI:2025zgx are correct, there might be a transition from $w<-1$ to $w>-1$ in the DE EoS.
In this section, we discuss if the (a little bit initial) condition determining if the standard phantom crossing (from $w>-1$ to $w<-1$) or the inverse phantom crossing (from $w<-1$ to $w>-1$) can be realised theoretically in the context of $F(R)$ gravity. Note that for future phantom crossing in $F(R)$ gravity, one can consult Bamba:2008hq . We show that the (inverse) phantom crossing in the $F(R)$ sector, which is described by the effective pressure and the effective energy density, can occur for almost any $F(R)$ gravity by adjusting the (initial) conditions.
@@
indicates that when the phantom crossing occurs, since we have $\rho=-p$, then $\dot{\rho}$ must vanish. Due to the fact that in the phantom Universe, $\rho<-p$, $\dot{\rho}$ and in the non-phantom Universe, $\rho>p$, there is a crossing from the non-phantom Universe to the phantom one if $\ddot{\rho}>0$ and there is an inverse crossing if $\ddot{\rho}<0$.
@@
Here $\rho_{\mathrm{matter(0)}}$ is a constant and $a$ is the scale factor, which is given in terms of the $e$-foldings $N$ as $a=\mathrm{e}^{N}$.
When the phantom crossing in the $f(R)$ sector happens at the cosmic time instance $t=t_{\mathrm{ph}}$, we assume $N$ can be expanded as,
@@
×$\,\times f^{\prime\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)$
@@
In the following, we consider the solution of the above equations and their physical meanings. Eq. (14) tells,
@@
By combining (13), (II), and (19), we find the following expression of $H_{2}$,
@@
Therefore, it depends on the detailed form of $f(R)$ if a crossing from the non-phantom to the phantom occurs, that is, $\rho_{{f(R)}\,(2)}>0$, or an inverse crossing, that is, $\rho_{{f(R)}\,(2)}<0$, occurs. This could indicate that the phantom crossing or the inverse phantom crossing occurs in a wide class of $F(R)$ gravity models, depending on the initial condition determining $H_{0}$. Furthermore, by combining (II) with (II), we can find the expression of $H_{3}$, although we do not write it explicitly because we do not use the expression of $H_{3}$ later.
@@
$f(R)=\alpha R^{2}\,.$ (23)
Here $\alpha$ is a constant. As is well-known, the $F(R)$ gravity includes a scalar mode. In the model (23), the square of the mass is given by $\frac{1}{\alpha}$. Therefore, in order to avoid the tachyon, we require $\alpha$ to be positive.
@@
For the standard phantom crossing to occur, the condition is given by
@@ III Apparent Phantom Crossing
There is an idea that instead of considering the transition from $w<-1$ to $w>-1$ of DE, the modification of the dark matter sector may solve the problem in the DESI observations Khoury:2025txd . If dark matter density decreases more slowly than $a^{-3}$, as usually predicted from energy conservation, it looks like there could be a phantom crossing in the DE sector because we are considering the total energy density, and we assume the usual $a^{-3}$ behaviour of the dark matter for the analysis of the DESI observations.
If the dark matter is non-relativistic, as in the cold dark matter, the energy density of the dark matter is given by the product $\rho_{\mathrm{DM}}=m_{\mathrm{DM}}n_{\mathrm{DM}}$ of the mass of the dark matter particle, $m_{\mathrm{DM}}$, and the number density of the dark matter particle, $n_{\mathrm{DM}}$. If there is neither creation nor annihilation of the dark matter particles, the number density $n_{\mathrm{DM}}$ is proportional to the inverse of the volume in the expanding Universe, $n_{\mathrm{DM}}\propto V^{-1}\propto a^{-3}$. If the mass $m_{\mathrm{DM}}$ is time dependent $m_{\mathrm{DM}}(t)$ or $m_{\mathrm{DM}}$ depends on the scalar curvature $R$, the behaviour of the dark matter energy density $\rho_{\mathrm{DM}}$ is different from $a^{-3}$. If the mass depends on the scalar curvature, the energy density also includes the derivative terms because the curvature includes the derivative, which makes the situation a little bit complicated.
In order to make the physical problem at hand well-defined, we assume that the dark matter is a Dirac spinor, whose action is given by
@@
[…]
The first term corresponds to the kinetic energy and the second term to the rest mass. As long as we consider non-relativistic particles, the first term can be neglected by comparing the second term. We should also note that $\bar{\psi}\psi\right>$ is nothing but the number density of dark matter $n_{\mathrm{DM}}$. Then we obtain,
@@
We assume that $\bar{\psi}\gamma_{i}D_{j}\psi\right>$ and $\left(D_{i}\bar{\psi}\right)\gamma_{j}\psi\right>$ are negligible for the non-relativistic dark matter. Then we find the energy density $\tilde{\rho}_{\mathrm{DM}}$ and the pressure $\tilde{p}_{\mathrm{DM}}$ as follows,
$\tilde{\rho}_{\mathrm{DM}}=$ $\,m(R)\left<\bar{\psi}\psi\right>+\left\{3\left(\dot{H}+H^{2}\right)+3H\frac{d}{dt}\right\}\left(m^{\prime}(R)\left<\bar{\psi}\psi\right>\right)$
@@
$\tilde{p}_{\mathrm{DM}}=$$\,\left\{-\left(\dot{H}+3H^{2}\right)+\frac{d^{2}}{dt^{2}}+2H\frac{d}{dt}\right\}\left(m^{\prime}(R)\left<\bar{\psi}\psi\right>\right)$
@@ IV Several Models that Exhibit Phantom Crossing: A Theoretical Account
where $\Lambda$ is a constant with dimension $[m]^{2}$ and $\beta$ is a dimensionless constant. When $R$ is large, $\beta\left(\frac{R}{2\Lambda}\right)^{\alpha}\gg 1$ if $\alpha>0$, and $\Lambda$ plays the role of the cosmological constant. This model is also proposed to solve the inverse phantom crossing. We choose $\frac{\Lambda}{\beta^{\frac{1}{\alpha}}}$ small enough compared with the present value of the scalar curvature. This makes $\Lambda$ give the DE in the present Universe. In order to also describe the inflation in the early Universe, we may add the term $\gamma R^{2}$ in the $F(R)$ of (35) with a constant $\gamma$.
Because $f(R)=-2\Lambda\left(1-\mathrm{e}^{-\beta\left(\frac{R}{2\Lambda}\right)^{\alpha}}\right)$, we find
@@
$\rho_{{f(R)}\,(2)}=$ $\,\frac{3}{\kappa^{2}}\left(-{H_{0}}^{2}H_{1}+{H_{1}}^{2}\right)-\frac{\Lambda^{2}\left(1-\mathrm{e}^{-\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}\right)\mathrm{e}^{\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}}{3\alpha\beta\kappa^{2}\left\{\left(\alpha-1\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{2\alpha-2}\right\}}$
@@
Then for example, when $\Lambda$ is small enough, $\rho_{{f(R)}\,(2)}\to\frac{3}{\kappa^{2}}\left(-{H_{0}}^{2}H_{1}+{H_{1}}^{2}\right)>0$. We should note $H_{1}$ is negative due to (19). Then, in the case that $\Lambda$ is small enough, the standard phantom crossing could occur. On the other hand, when $\Lambda$ is large enough, we find $\rho_{{f(R)}\,(2)}\to-\frac{\Lambda^{2}}{3\beta^{2}\kappa^{2}}\left(1-\mathrm{e}^{-\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)}\right)<0$, and as long as $-H_{1}<2{H_{0}}^{2}$, the inverse phantom crossing could occur.
@@
+$\,+\left\{3\left(\dot{H}+H^{2}\right)+3H\frac{d}{dt}\right\}\left\{\frac{1}{2\kappa^{2}}\left(1+f^{\prime}(R)\right)-m^{\prime}(R)n_{\mathrm{DM}}\right\}\,.$ (39)
By using the $e$-foldings $N$ defined by $a=\mathrm{e}^{N}$ by assuming $a=1$ in the present Universe, we find $n_{\mathrm{DM}}=n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}$, where $n_{\mathrm{DM\,(0)}}$ is the number density of the dark matter particles in the present Universe. In terms of $e$-foldings $N$, Eq. (IV) can be rewritten as,
@@
+$\,+\left\{3\left(HH^{\prime}+H^{2}\right)+3H^{2}\frac{d}{dN}\right\}\left\{\frac{1}{2\kappa^{2}}\left(1+f^{\prime}(R)\right)-m^{\prime}(R)n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}\right\}\,,$ (40)
@@
Using the expression of $N=N(R)$, we find $m(R)$ as a function of $R$. Then by the definition of $\Psi$, we find,
@@
Then we find $f(R)$ as a function of $R$. First, we consider a very trivial example,
@@
with constants $H_{0}$ and $\Phi_{0}$ to check the consistency in the above formulation. The spacetime is, of course, the de Sitter spacetime. Then Eqs. (IV), (43), and (44) tell,
@@
Then we obtained the model with a cosmological constant, where $f(R)$ is a constant and there is no matter $m(R)=0$, as expected. Anyway, we find the above formulation using Eqs. (IV), (43), and (44) may work. Second, as a less non-trivial example, we consider the case that
@@
Therefore, we obtain Einstein’s gravity with a cosmological constant consistently, although the gravitational coupling $\kappa$ is effectively modified as $\kappa^{2}\to\frac{n_{\mathrm{DM\,(0)}}}{\Phi_{0}}$ because $F(R)=R+f(R)=\frac{2\kappa^{2}\Phi_{0}}{n_{\mathrm{DM\,(0)}}}R+6\kappa^{2}{H_{0}}^{2}\Phi_{0}\left(-1+\frac{4}{n_{\mathrm{DM\,(0)}}}\right)$. We now consider a more realistic and non-trivial model. As we found, the inverse phantom crossing occurs when the first derivative of the effective DE density $\rho^{\mathrm{eff}}_{\mathrm{DE}}$ vanishes and its second derivative is negative,
@@
The $e$-foldings $N$ is a monotonically increasing function of the cosmic time $t$, and the above condition (49) can be rewritten in terms of $N$, as follows,

### v2

@@ Abstract
In this work, we shall consider how a dynamical oscillating and phantom crossing dark energy era can be realized in the context of $F(R)$ gravity. We approach the topic from a theoretical standpoint considering all the conditions that may lead to a consistent phantom crossing behavior and separately how the $F(R)$ gravity context may realize oscillating dark energy era. Apart from our qualitative considerations, we study in a quantitative way two $F(R)$ gravity dark energy models which are viable cosmologically and also exhibit simultaneously phantom crossing behavior and also oscillating dark energy. We consider these models by solving numerically the field equations using appropriate statefinder parameters engineered for dark energy studies. As we show, $F(R)$ provides a natural extension of Einstein’s general relativity which can naturally realize a transition from a phantom era to a quintessential era, a feature supported by recent observational data, without resorting to phantom scalar fields to realize the phantom evolution.
@@ I Introduction
The observation of the late-time acceleration of the Universe, known as dark energy (DE) era is one of the most unexpected features of our Universe. In the context of simple general relativity, the DE era can be generated by quintessence scalar fields, however the latest observations of the Dark Energy Spectroscopic Instrument (DESI) DESI:2024mwx ; DESI:2025zgx indicate that the DE was in the phantom regime in the near past. This feature can be described in the context of general relativity only by using phantom scalar fields, which is a rather unattractive feature. Also and more importantly, the DESI data indicate a transition from a phantom to a quintessential regime for DE. This is questionable if it can be realized in the context of general relativity. There are many ways to produce results compatible with the DESI data, for example OdintsovSGS:2024 ; Dai:2020rfo ; He:2020zns ; Nakai:2020oit ; DiValentino:2020naf ; Agrawal:2019dlm ; Ye:2020btb ; Vagnozzi:2021tjv ; Desmond:2019ygn ; Hogas:2023pjz ; OColgain:2018czj ; Vagnozzi:2019ezj ; Krishnan:2020obg ; Colgain:2019joh ; Vagnozzi:2021gjh ; Lee:2022cyh ; Krishnan:2021dyb ; Ye:2021iwa ; Ye:2022afu ; Verde:2019ivm ; Menci:2024rbq ; Adil:2023ara ; Reeves:2022aoi ; Ferlito:2022mok ; Vagnozzi:2021quy ; DiValentino:2020evt ; Sabogal:2024yha ; Giare:2024smz ; DiValentino:2025sru ; Odintsov:2025kyw ; vanderWesthuizen:2025iam ; Paliathanasis:2025dcr ; Odintsov:2025jfq ; Kessler:2025kju and among these frameworks, $F(R)$ gravity reviews1 ; reviews2 ; reviews3 ; reviews4 ; Nojiri:2003ft ; Capozziello:2005ku ; Hwang:2001pu ; Song:2006ej ; Faulkner:2006ub ; Olmo:2006eh ; Sawicki:2007tf ; Faraoni:2007yn ; Carloni:2007yv ; Nojiri:2007as ; Deruelle:2007pt ; Appleby:2008tv ; Dunsby:2010wg ; Oikonomou:2025qub enjoys an elevated role. There are recent studies which discuss the ability of $F(R)$ gravity to be compatible with the DESI data Odintsov:2025jfq . Another more fascinating behavior of DE based on the DESI data was demonstrated in Ref. Kessler:2025kju , which indicates that the DE might not be only dynamical and evolving from phantom to quintessence, but it also shows strong indications of an oscillating equation of state (EoS). In this work we aim to investigate whether a phantom involving and oscillating DE may be realized by $F(R)$ gravity in general. We approach the problem theoretically, finding the conditions which may allow such a phenomenological behavior to be produced and also we present several phenomenological models from a theoretical point of view. In addition, we study numerically the DE era of some explicit $F(R)$ gravity models which can realize a viable Planck compatible Planck:2018vyg and in some cases DESI compatible late-time behavior. We also demonstrate that these models produce an oscillating DE era and also allow for phantom to quintessence transitions.
@@
This paper is organized as follows: In section II we discuss the inverse-phantom prossing in $F(R)$ gravity from a theoretical viewpoint. The analysis is continued in section III where we discuss the apparent phantom crossing. In section IV, the theoretical study of several models that exhibit phantom crossing is considered. The oscillating Hubble rate is considered in section V, while in section VI the effects of a scalaron as dark matter with curvature-dependent mass on the DE era are considered. In section VII, we discuss the features of a realistic model while in section VIII we numerically study two explicit examples of oscillating $F(R)$ gravity DE with phantom crossing. Finally, the conclusions follow at the end of the article.
@@ II Inverse-Phantom Crossing in $F(R)$ gravity
$\centering\mathcal{S}=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}F(R)\,$,\@add@centering (1)
@@
Let $w$ be the DE equation of state parameter, that is, the ratio of the pressure of the DE fluid with respect to the energy density of the DE fluid. If the recent DESI observations DESI:2024mwx ; DESI:2025zgx are correct, there might be a transition from w<1w<-1 to w>1w>-1 in the DE EoS.
In this section, we discuss if the initial condition determining if the standard phantom crossing (from w>1w>-1 to w<1w<-1) or the inverse phantom crossing (from w<1w<-1 to w>1w>-1) can be realized theoretically in the context of $F(R)$ gravity. Therefore, we show that the phantom crossing (or inverse phantom crossing) in the $F(R)$ sector (which is described by the effective pressure and the effective energy density) can occur for almost any $F(R)$ gravity by adjusting the initial conditions.
@@
which indicates that when the phantom crossing occurs, since we have $\rho=-p$, then $\dot{\rho}$ must vanish. Due to the fact that in the phantom Universe, $\rho<-p$, $\dot{\rho}$ and in the non-phantom Universe, $\rho>p$, there is a crossing from the non-phantom Universe to the phantom one if $\ddot{\rho}>0$ and there is an inverse crossing if $\ddot{\rho}<0$.
@@
Here $\rho_{\mathrm{matter(0)}}$ is a constant and $a$ is the scale factor, which is given in terms of the $e$-foldings number $N$ as $a=\mathrm{e}^{N}$.
When the phantom crossing in the $f(R)$ sector happens at the cosmic time instance $t=t_{\mathrm{ph}}$, we can assume that $N$ can be expanded as follows,
@@
$\,\times f^{\prime\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)$
@@
In the following, we consider the solution of the above equations and their physical meanings. Eq. (14) indicates that,
@@
By combining (13), (II), and (19), we find the following expression for $H_{2}$,
@@
Therefore, it depends on the detailed form of $f(R)$ if a crossing from the non-phantom to the phantom occurs, that is, $\rho_{{f(R)}\,(2)}>0$, or an inverse crossing, that is, $\rho_{{f(R)}\,(2)}<0$, occurs. This could indicate that the phantom crossing or the inverse phantom crossing occurs in a wide class of $F(R)$ gravity models, depending on the initial condition determining $H_{0}$. Furthermore, by combining (II) with (II), we can find the expression of $H_{3}$ although we do not write it explicitly because we do not use the expression of $H_{3}$ later.
@@
$f(R)=\alpha R^{2}\,$, (23)
where $\alpha$ is a constant. As is well-known, the $F(R)$ gravity includes a scalar mode. In the model (23), the square of the mass is given by $\frac{1}{\alpha}$. Therefore, in order to avoid the tachyon, we require $\alpha$ to be positive.
@@
For the standard phantom crossing to occur, the condition is given by,
@@ III Apparent Phantom Crossing
There is an idea that instead of considering the transition from w<1w<-1 to w>1w>-1 of DE, the modification of the dark matter sector may solve the problem in the DESI observations Khoury:2025txd . If dark matter density decreases more slowly than $a^{-3}$, as usually predicted from energy conservation, it looks like there could be a phantom crossing in the DE sector because we are considering the total energy density, and we assume the usual $a^{-3}$ behavior of the dark matter for the analysis of the DESI observations.
If the dark matter is non-relativistic, as in the cold dark matter, the energy density of the dark matter is given by the product $\rho_{\mathrm{DM}}=m_{\mathrm{DM}}n_{\mathrm{DM}}$ of the mass of the dark matter particle, $m_{\mathrm{DM}}$, and the number density of the dark matter particle, $n_{\mathrm{DM}}$. If there is neither creation nor annihilation of the dark matter particles, the number density $n_{\mathrm{DM}}$ is proportional to the inverse of the volume in the expanding Universe, $n_{\mathrm{DM}}\propto V^{-1}\propto a^{-3}$. If the mass $m_{\mathrm{DM}}$ is time dependent $m_{\mathrm{DM}}(t)$ or $m_{\mathrm{DM}}$ depends on the scalar curvature $R$, the behavior of the dark matter energy density $\rho_{\mathrm{DM}}$ is different from $a^{-3}$. If the mass depends on the scalar curvature, the energy density also includes the derivative terms because the curvature includes the derivative, which makes the situation somewhat complicated.
In order to make the physical problem at hand well-defined, we assume that the dark matter is a Dirac spinor, whose action is given by,
@@
$\rho_{\mathrm{DM}}=\left<\mathcal{H}\right>=-i\left<\bar{\psi}\gamma^{i}\partial_{i}\psi\right>+m(R)\left<\bar{\psi}\psi\right$>\,. (29)
The first term corresponds to the kinetic energy and the second term to the rest mass. As long as we consider non-relativistic particles, the first term can be neglected by comparing the second term. We should also note that $\left<\bar{\psi}\psi\right$> is nothing but the number density of dark matter $n_{\mathrm{DM}}$. Then we obtain,
@@
We assume that $\left<\bar{\psi}\gamma_{i}D_{j}\psi\right$> and $\left<\left(D_{i}\bar{\psi}\right)\gamma_{j}\psi\right$> are negligible for the non-relativistic dark matter. Then we find the energy density $\tilde{\rho}_{\mathrm{DM}}$ and the pressure $\tilde{p}_{\mathrm{DM}}$ as follows,
[…]
@@
[…]
@@ IV Several Models that Exhibit Phantom Crossing: A Theoretical Account
where $\Lambda$ is a constant with dimension $m]^{2}$ and $\beta$ is a dimensionless constant. When $R$ is large, $\beta\left(\frac{R}{2\Lambda}\right)^{\alpha}\gg 1$ if $\alpha>0$, and $\Lambda$ plays the role of the cosmological constant. This model is also proposed to solve the inverse phantom crossing. We choose $\frac{\Lambda}{\beta^{\frac{1}{\alpha}}}$ small enough compared with the present value of the scalar curvature. This makes $\Lambda$ to realize a DE era in the present Universe. In order to also describe the inflation in the early Universe, we may add the term $\gamma R^{2}$ in the $F(R)$ of (35) with a constant $\gamma$.
Because $f(R)=-2\Lambda\left(1-\mathrm{e}^{-\beta\left(\frac{R}{2\Lambda}\right)^{\alpha}}\right)$, we find,
@@
$\rho_{{f(R)}\,(2)}=$ $\,\frac{3}{\kappa^{2}}\left(-{H_{0}}^{2}H_{1}+{H_{1}}^{2}\right)-\frac{2\Lambda\left(1-\mathrm{e}^{-\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}\right)\mathrm{e}^{\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}}{12\kappa^{2}\frac{\alpha\beta}{2\Lambda}\left\{\left(\alpha-1\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{2\alpha-2}\right\}}$
$\,+\frac{\left({H_{0}}^{2}+H_{1}\right)\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-1}}{2\kappa^{2}\frac{\alpha\beta}{2\Lambda}\left\{\left(\alpha-1\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{2\alpha-2}\right\}}$
$\,+\frac{\left(3{H_{0}}^{2}+H_{1}\right)\mathrm{e}^{\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}}{6\kappa^{2}\frac{\alpha\beta}{2\Lambda}\left\{\left(\alpha-1\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{2\alpha-2}\right\}}$
$=$ $\,\frac{3}{\kappa^{2}}\left(-{H_{0}}^{2}H_{1}+{H_{1}}^{2}\right)-\frac{\Lambda^{2}\left(1-\mathrm{e}^{-\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}\right)\mathrm{e}^{\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}}{3\alpha\beta\kappa^{2}\left\{\left(\alpha-1\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{2\alpha-2}\right\}}$
@@
Then for example, when $\Lambda$ is small enough, $\rho_{{f(R)}\,(2)}\to\frac{3}{\kappa^{2}}\left(-{H_{0}}^{2}H_{1}+{H_{1}}^{2}\right)>0$. We should note that $H_{1}$ is negative due to (19). Then in the case that $\Lambda$ is small enough, the standard phantom crossing could occur. On the other hand, when $\Lambda$ is large enough, we find $\rho_{{f(R)}\,(2)}\to-\frac{\Lambda^{2}}{3\beta^{2}\kappa^{2}}\left(1-\mathrm{e}^{-\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)}\right)<0$, and as long as $H_{1}<2{H_{0}}^{2}$, the inverse phantom crossing could occur.
@@
$\,+\left\{3\left(\dot{H}+H^{2}\right)+3H\frac{d}{dt}\right\}\left\{\frac{1}{2\kappa^{2}}\left(1+f^{\prime}(R)\right)-m^{\prime}(R)n_{\mathrm{DM}}\right\}\,.$ (39)
By using the $e$-foldings NN defined by $a=\mathrm{e}^{N}$ by assuming $a=1$ in the present Universe, we find $n_{\mathrm{DM}}=n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}$, where $n_{\mathrm{DM\,(0)}}$ is the number density of the dark matter particles in the present Universe. In terms of the $e$-foldings NN, Eq. (IV) can be rewritten as,
@@
$\,+\left\{3\left(HH^{\prime}+H^{2}\right)+3H^{2}\frac{d}{dN}\right\}\left\{\frac{1}{2\kappa^{2}}\left(1+f^{\prime}(R)\right)-m^{\prime}(R)n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}\right\}\,$, (40)
@@
Using the expression of $N=N(R)$, we find $m(R)$ as a function of $R$. Then by using the definition of $\Psi$, we find,
@@
Then we can find $f(R)$ as a function of $R$. First, we consider a very trivial example,
@@
with constants $H_{0}$ and $\Phi_{0}$ to check the consistency in the above formulation. The spacetime is, of course, the de Sitter spacetime. Then Eqs. (IV), (43), and (44) indicate that,
@@
Hence we obtained the model with a cosmological constant, where $f(R)$ is a constant and there is no matter $m(R)=0$, as expected. We can find the above formulation using Eqs. (IV), (43), and (44) may work. Second, as a less non-trivial example, we consider the case that,
@@
Therefore, we obtain Einstein’s gravity with a cosmological constant consistently although the gravitational coupling $\kappa$ is effectively modified as $\kappa^{2}\to\frac{n_{\mathrm{DM\,(0)}}}{\Phi_{0}}$ because $F(R)=R+f(R)=\frac{2\kappa^{2}\Phi_{0}}{n_{\mathrm{DM\,(0)}}}R+6\kappa^{2}{H_{0}}^{2}\Phi_{0}\left(-1+\frac{4}{n_{\mathrm{DM\,(0)}}}\right)$. We now consider a more realistic and non-trivial model. As we found, the inverse phantom crossing occurs when the first derivative of the effective DE density $\rho^{\mathrm{eff}}_{\mathrm{DE}}$ vanishes and its second derivative is negative,
@@
The $e$-foldings NN is a monotonically increasing function of the cosmic time $t$, and the above condition (49) can be rewritten in terms of $N$, as follows,